OpenStudy (anonymous):

f(x) =2x - 33 for x=9???

OpenStudy (teddyiswatshecallsme):

What are the instructions?

OpenStudy (anonymous):

evaluate each function for the given value of x, and write the input x and output f(x) as an ordered pair @Teddyiswatshecallsme

OpenStudy (teddyiswatshecallsme):

ooh i ahven't gotten to functions yet, sorry. ;-; I CAN'T HELP... Ifeel so useles.s.

OpenStudy (anonymous):

its okay thank you!

OpenStudy (johnweldon1993):

we have a function \[\large f(x) = 2x - 33\] and we are told to figure out what that function will equal...when x = 9 So to find that out...all we do is replace every 'x' we see in the original equation...with 9 \[\large f(9) = 2(9) - 33\] \[\large f(9) = 18 - 33\] \[\large f(9) = ?\]

OpenStudy (anonymous):

f(9) = -15

OpenStudy (johnweldon1993):

Perfect!

OpenStudy (anonymous):

it says write the input and output values as an ordered pair?

OpenStudy (anonymous):

oh and could you help me with one more please

OpenStudy (johnweldon1993):

Okay, so the input is the 'x' value....and the output is what the function equals with that input value.. So what did we input? we did x = 9 what did that output? it outputted y = -15 So the ordered pair would be (x,y) = (9 , -15)

OpenStudy (anonymous):

ohhhh okay okay awesome

OpenStudy (johnweldon1993):

And of course I'll help you out with another :P

OpenStudy (anonymous):

yayyy

OpenStudy (anonymous):

f(x) = -2x +1/3 for x = -5

OpenStudy (johnweldon1993):

Alright so same thing here (gonna make you do it this time :P) we have \[\large f(x) = -2x + \frac{1}{3}\] and they want to know what that equals if you make 'x' = -5

OpenStudy (anonymous):

my ordered pair i got was (-5, 11/3)

OpenStudy (anonymous):

its actually \[-2x + 1\div3\]

OpenStudy (anonymous):

not just 1/3 its all over the 3

OpenStudy (johnweldon1993):

Oh its the full thing \[\large \frac{-2x + 1}{3}\] Then yes your ordered pair is correct!

OpenStudy (anonymous):

sorry can you check just one more for me and then i should be good ahah

OpenStudy (johnweldon1993):

lol of course :)

OpenStudy (anonymous):

f(x) - 2/9x for -9/2 for x =9 and i got (9, -9) for my ordered pair

OpenStudy (anonymous):

@johnweldon1993